OpenStudy (anonymous):

Find the derivative of the function f(x) = 3 sqrt x + 2e^x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

ok thanks

myininaya (myininaya):

f(x)=3x^(1/2)+2e^x f'(x)=3*1/2*x^(1/2-1)+2e^x =3/2*x^(-1/2)+2e^x

OpenStudy (anonymous):

what happens the to the 2e^x

myininaya (myininaya):

(e^x)'=e^x

myininaya (myininaya):

the derivative is itself

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

thats just the rules?

myininaya (myininaya):

e^x is an increasing function (e^x)' is going to be above the x-axis (but e^x is never negative anyways) so can you see how (e^x)'=e^x possibly now?

OpenStudy (anonymous):

yes thank you

myininaya (myininaya):

and yes they did make it a rule but it can be proven

OpenStudy (anonymous):

do you have to ask a certain amount of questions to give a person a good answer because am not able to see it the lady said it was suppose to be in the right hand corner

OpenStudy (anonymous):

myininaya can you help on the problem i posted please?

OpenStudy (anonymous):

its trignometric substitution